OpenStudy (anonymous):

write a quadratic equation in the variable x having the given numbers as solutions type the equation in standard form, ax^2+bx+c=0

OpenStudy (anonymous):

ax^2+bx+c=0 in standard form?

OpenStudy (anonymous):

thatswhat the question asked

OpenStudy (sriram):

what are the given numbers?

OpenStudy (anonymous):

quadratic formula

OpenStudy (anonymous):

the question is writen as it was given to me

OpenStudy (anonymous):

i guess we need the given numbers. i am not trying to give you a hard time, but make sure you understand the question. there is something missing here

OpenStudy (anonymous):

x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}

OpenStudy (anonymous):

it asks for a quadratic equation whose solutions are given. you need to know the numbers that are given, otherwise this makes no sense

OpenStudy (anonymous):

quadratic formula

OpenStudy (anonymous):

say the "given numbers" are 2 and 5. then you write \[(x-2)(x-5)=0\] and if you like multiply out to get \[x^2-7x+10=0\] that is a quadratic equation with the "given" zeros of 2 and 5

OpenStudy (anonymous):

ummm...

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ax^2+bx+c=0 thats all thats given

OpenStudy (anonymous):

it is not "quadratic formula" it is "find the equation with the GIVEN zeros"

OpenStudy (anonymous):

ok satellite

OpenStudy (anonymous):

solution 2, only solution sorry the paper was cut

OpenStudy (anonymous):

@ panther i have seen these questions repeatedly. either there is some mistake in the question, or somewhere there are two numbers given to be the zeros. trust me on this

OpenStudy (anonymous):

ok now we got somewhere. the answer is \[(x-2)^2=0\]

OpenStudy (anonymous):

i have given you all there was sorry i didnt see it at first

OpenStudy (anonymous):

you have to multiply out to get \[x^2-4x+4=0\] and that is your answer

OpenStudy (anonymous):

ty